Project

General

Profile

NDOS Commissioning Meeting 2011/03/03

Attendance: Jon, Andrew, Nate, Gavin, Jarek, Minerba, Mark, Mat, Gary, Leon, Alena, Susan, Athans, Rick K., Dnis, Tricia, Peter, Chad

Agenda:

  1. Channel Recovery Plans
  2. DAQ Update & Plans
  3. DCS Status & Plans
  4. Shifters' report
  5. AOB

Channel Recovery (Mat):

  • Jon and Mat recovered some more channels today. Two FEBs were found to have HV ribbon cables disconnected.
  • All "red" APDs have been removed, only 8 more need to be replaced.
  • Investigated 3 FEBs that were misbehaving. One FEB had ribbon cables partially inserted, swapped 2 others. One of the swapped looked good afterward, the other is not clear yet if it helped.
  • Performed a cablectomy and this seemed to improve the readout for all channels.
  • Looked at most recent DSO scan, 62/63 of recovered APD/FEBs look good since they were "fixed".
  • Denis reported that some channels behave poorly immediately after they were cooled, but eventually get better.
  • Jon and Mat also swapped TEC boards on FEB# 18 and #19 on DCM 1-2, as requested by Gerard and Denis.

DSO Studies (Mark Mathis):

DAQ Software:

  • Peter: several things have happened this week:
    • DAQ has moved to new version of underlying messaging system in an attempt to make the system more robust.
    • hopefully this will make stopping and restarting the messaging system between runs unnecessary (tests are in progress)
    • small change in DCM in bus arbitration that increases capability of memory transfer (10% decrease in CPU usage)
    • Peter has demonstrated that configured threshold matches observed threshold
    • work continues on increasing DCM bandwidth

DCS:

  • Chad: asked Leon if RTD purchase is approved.
  • Leon: concerned about soldering proposal
  • RTDs require silver solder... Feynman has experience doing this
  • Chad has placed an order for all of the infrastructure parts required to mount the RTDs.
  • Athans: managed to test code on DCM on teststand and is seeing reasonable values.

Any other business:

  • Andrew: Would like to schedule some downtime to remap network resources on both detector and LCC. Estimate ~2 hours at each site. Tentatively scheduled for Monday morning.
  • Mark: efforts to search for MB events in detector is ramping up (Minerba + Chad). Booster is getting ~2x amount of spills because NuMI is down. MC is being developed for MB events.